13. A population has a standard deviation of 16. If a sample of size 64 is...

13.

A population has a standard deviation of 16. If a sample of size 64 is selected from this population, what is the probability that the sample mean will be within ±2 of the population mean?

Select one:

a. 0.6826

b. -0.6826

c. 0.3413

d. Since the mean is not given, there is no answer to this question.
Answer #1

Standard deviation, = 16

Sample size, n = 64

Standard error =

=

= 2

P(sample mean will be within ±2 of the population mean) = P(Z < 2/standard error) - P(Z < -2/standard error)

= P(Z < 2/2) - P(Z < -2/2)

= P(Z < 1) - P(Z < -1)

= 0.8413 - 0.1587

= 0.6826